OpenStudy (anonymous):

Calculate fog and gof : f(x)=e^x^2 g(x)=e^-x^2

OpenStudy (helder_edwin):

it is \[\large f(x)=e^{x^2} \] and \[\large g(x)=e^{-x^2} \] right?

OpenStudy (helder_edwin):

just use the definition \[\large (f\circ g)(x)=f(g(x)). \]

OpenStudy (anonymous):

??????????

OpenStudy (ikram002p):

wat r u stuck in ??

OpenStudy (anonymous):

؟

OpenStudy (helder_edwin):

in the first case, the idea is to replace g(x) in every x of f(x)

OpenStudy (ikram002p):

i mean which step you dont understand ??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the finale step

OpenStudy (ikram002p):

fog =f(g(x)) ??

OpenStudy (anonymous):

fog(x)=e^e^(-x^2)^2 ??

OpenStudy (ikram002p):

would u like me to explain it to u in arabic ?? " i suppose u know arabic right ??"

OpenStudy (anonymous):

نعممممممممممممممممممممم

OpenStudy (ikram002p):

fog معناها f بعد g يعني لازم تشوف بالأول الاقتران g وتطبقو على f حتى الآن جيد ؟؟

OpenStudy (anonymous):

اوك

OpenStudy (ikram002p):

طيب g(x)=e^(-x)^2=e^(-2x) ماشي ؟؟

OpenStudy (anonymous):

ماشي

OpenStudy (ikram002p):

اوك هسا بدك تشوف f(g(x)) عوض بدل g(x) القيمة اللي طلعت عندك يعني بدك تشوف f(e^(-2x))

OpenStudy (anonymous):

اكتبي كل الاجابات بليز

OpenStudy (ikram002p):

اوكي ما انا بعمل خطوة بخطوة

OpenStudy (ikram002p):

لحد هون فاهم ؟؟

OpenStudy (ikram002p):

f(x)=e^x^2 =e^(2x) بس انك بدك توجد fog=f(g(x)) يعني مكان كل x بالاقتران عوض g(x) f(g(x))=e^(2g(x)) =e^(2(e^(-2x))) =e^(-4xe)=-4x

OpenStudy (ikram002p):